# Exploring the Dynamics of General Relativistic Binary-Single and Binary-Binary Encounters of Black Holes

### v1

@@ IV.3 Almost Head-on Collision After a Strong Binary-Single Scattering (BSS-03)
Figure 13: Left: The real part (blue), the imaginary part (orange) and the absolute value (green) of $R\Psi_{4}$ for the system BSS-03 extracted at the radius $R=200$ M and on the level $L=2$ in the direction $\theta=0.9,\phi=3.5$. $R\Psi_{4}$ is reconstructed from all modes from $l=2$ to $l=4$. Right: Gravitational-wave luminosity extracted at the same radius and level. The colored dashed lines mark the maxima in gravitational-wave luminosity and correspond to the crosses in Figure 11 and 12.
@@
Figure 13 shows $R\Psi_{4}$ in a specific direction ($\theta=0.9,\phi=3.5$) as well as the gravitational-wave luminosity, both extracted at the radius $R=200$ M and on the level $L=2$. At the beginning of the simulation, one can see small oscillations in the real and imaginary parts of $R\Psi_{4}$ due to the binary inspiral. The insprial is interrupted by three distinct peaks, and the times at which the peak maxima occur are indicated by differently colored dashed lines. The first two gravitational-wave bursts are primarily sent in the positive and negative $z$-directions, whereas the third burst is primarily directed perpendicular to the $z$-axis. Because of this, we choose a direction for $R\Psi_{4}$ in which all the peaks that are visible in the gravitational-wave luminosity also clearly appear in the plot of $R\Psi_{4}$. For this system, one can also connect the peaks in the gravitational-wave emission to specific events in the binary-single encounter. We again illustrate this connection by adding crosses to the puncture trajectory plots in Figures 11 and 12 at the retarded times $t_{R}=t-R$, which approximately correspond to the moments in time $t$ marked by the dashed lines in Figure 13. One can see that the first peak (colored red) corresponds to the moment when the black holes in the binary get rapidly accelerated out of the orbital plane. This is especially visible in the right plot in Figure 12.
@@
Figure 15: The real part (blue), imaginary part (orange), and the absolute value (green) of $R\Psi_{4}$ for the system BSS-04, extracted at the radius $R=140$ M and on the level $L=2$ in the direction $\theta=0,\phi=0$ (which corresponds to the $z$-axis).
@@ IV.4 Highly Eccentric Merger After a Strong Binary-Single Scattering (BSS-04)
Our fourth system moves in the $xy$-plane with the normal vector of the binary’s orbital plane pointing in the $z$-direction. For this system we pick $X=50$ M, $b=40$ M and $|\vec{p}|=0.057282195$ M. The black hole trajectories of the entire system and of the binary in the center-of-mass frame are depicted in Figure 14. Figure 15 shows $R\Psi_{4}$ in a specific direction ($\theta=0,\phi=0$, which corresponds to the $z$-axis, i.e., the direction of the normal vector of the binary’s orbital plane), extracted at the radius $R=140$ M and on the level $L=2$. One can see that the single black hole induces a significantly accelerated merger with a high eccentricity. In the comparison with the 2.5PN approximation, one can observe that the dynamics look quite similar up to a certain point, although distorted in the direction of motion as in the other examples. The scattering angle appears noticeably larger in the post-Newtonian approximation, which is contrary to what has been observed for the strong-field scattering of individual black holes in [71], where the scattering angles of the higher-order post-Newtonian approximations are always smaller than those of the fully relativistic cases.
@@
Figure 16: Top left: Puncture trajectories of the black holes of BSS-05 (colored blue for the black holes in the binary and orange for the single black hole) with a comparison to the 2.5PN approximation. Top right: Gravitational-wave luminosity extracted at the radius $R=160$ M and on the level $L=2$. Bottom: The real part (blue), the imaginary part (orange), and the absolute value (green) of $R\Psi_{4}$, extracted at the same radius and level and in the direction $\theta=2.1,\phi=0$.
@@ IV.5 Exchange and Subsequent Merger After a Strong Binary-Single Scattering (BSS-05)
Figure 17: Left: Gravitational-wave luminosity extracted at the radius $R=300$ M and on the level $L=2$. Right: The real part (blue), the imaginary part (orange), and the absolute value (green) of the 22-mode of $R\Psi_{4}$, extracted at the same radius and level and in the direction $\theta=0.5,\phi=4.9$.
@@
The plot of $R\Psi_{4}$ in Figure 16 shows the regular binary signal, interrupted by a small bump and a subsequent decay of the signal, which is then followed by a stronger burst of radiation. This can also be seen in the plot of $dE/dt$ where the bump at $t\approx 600$ M is relatively small. Looking at the system at the retarded time $t_{R}=t-R$, one can conclude that the small bump arises from the single black hole slowing down one black hole and accelerating the other one, whereas the burst at the end originates from the merger of the single black hole with one of the binary members. It might be the case that the existence of a bump followed by a binary signal with a different frequency, amplitude, and waveform (or a prompt merger as in this case) is a characteristic feature of an exchange since it requires the deceleration of one binary member’s circular motion and a strong acceleration of the other binary member, leading to a moderate second derivative of the binary’s mass quadrupole moment.
@@ IV.6 Strong Binary-Binary Scattering (BBS-01)
The plot of $dE/dt$ in Figure 17 shows two distinct and irregularly shaped peaks, as well as a smaller constant emission around the time of closest approach (red dashed line). The first peak (colored magenta) likely corresponds to the strong acceleration of both binaries and/or the acceleration of individual binary members. At the time of the first peak, the black holes in the blue-colored binary get accelerated outwards and reach their widest separation throughout the simulation, while at the same time the black-hole distance in the eccentric orange binary reaches a local minimum. The second peak (colored turquoise) is associated with the merger of the orange-colored binary. One can see that the maximum of the absolute value of $R\Psi_{4}$ occurs earlier than in the gravitational-wave luminosity. This is because the merger sends two bursts in opposite directions. One burst reaches the extraction sphere before the time marked by the turquoise line, while the other one hits it afterward. The plot depicting $R\Psi_{4}$ also shows the initial signal of both binaries. Due to the different phases, the amplitude gets slightly modulated. The final part of the signal comes from the gravitational-wave emission of the remaining eccentric binary. In the center-of-mass systems of each binary, the black holes get accelerated out of the initial orbital plane during the encounter, and the orbital plane of the binary with blue trajectories gets permanently inclined by around 10 degrees relative to its initial orbital plane.
@@ V Summary and Conclusions
In addition to the previously mentioned difficulties, the extraction of the gravitational-wave signal of binary-single and binary-binary encounters is also challenging due to the fact that the binary is not at rest at the center of the extraction sphere. There are works that deal with the signal extraction for binaries whose center of mass is displaced from the origin or moves at a constant velocity [75, 76, 77], but in our case the binary is accelerating and rotating, which leads to mode mixing and spurious oscillations of $|R\Psi_{4}|$. When looking at the gravitational-wave luminosity that takes into account all modes, these oscillations are significantly reduced, but some peaks still show irregular shapes, and it is unclear to what extent this is a physical effect or a result of our method of gravitational-wave signal extraction.
@@ Acknowledgements
Animated videos of some of the plots presented in this paper can be found at https://doi.org/10.5281/zenodo.15342415 [90]. They give valuable insights that are not obvious from the plots. The animated version of Figure 9 was created with rose [91]. The code for integrating the three-body post-Newtonian equations of motion up to 2.5PN is available at https://github.com/fmheinze/pn-nbody. The data analyzed in this work are available upon request from the corresponding author.
Appendix A Extra Material
Table 4 provides the full list of parameters used for all the systems in the main text. In Figure 20 we show the 22-mode and the 44-mode of $R\Psi_{4}$ for a quasi-circular binary that is moving uniformly with respect to the origin of the gravitational-wave extraction sphere as a simple example of mode mixing and the associated spurious oscillations that are discussed in the main text.
@@ Appendix A Extra Material
Figure 20: The real part (blue), the imaginary part (orange), and the absolute value (green) of the 22-mode (top panel) and the 44-mode (bottom panel) of $R\Psi_{4}$ for a quasi-circular black-hole binary moving with a momentum $|\vec{p}|=0.05$ M in the positive $x$-direction, away from the coordinate origin. The signal is extracted at the radius $R=300$ M and on the level $L=3$.

### v2

@@ IV.3 Almost Head-on Collision After a Strong Binary-Single Scattering (BSS-03)
Figure 13: Left: The real part (blue), the imaginary part (orange) and the absolute value (green) of $RM\Psi_{4}$ for the system BSS-03 extracted at the radius $R=200M$ and on the level $L=2$ in the direction $\theta=0.9,\phi=3.5$. $RM\Psi_{4}$ is reconstructed from all modes from $l=2$ to $l=4$. Right: Gravitational-wave luminosity extracted at the same radius and level. The colored dashed lines mark the maxima in gravitational-wave luminosity and correspond to the crosses in Figure 11 and 12.
@@
Figure 13 shows $RM\Psi_{4}$ in a specific direction ($\theta=0.9,\phi=3.5$) as well as the gravitational-wave luminosity, both extracted at the radius $R=200M$ and on the level $L=2$. At the beginning of the simulation, one can see small oscillations in the real and imaginary parts of $RM\Psi_{4}$ due to the binary inspiral. The insprial is interrupted by three distinct peaks, and the times at which the peak maxima occur are indicated by differently colored dashed lines. The first two gravitational-wave bursts are primarily sent in the positive and negative $z$-directions, whereas the third burst is primarily directed perpendicular to the $z$-axis. Because of this, we choose a direction for $RM\Psi_{4}$ in which all the peaks that are visible in the gravitational-wave luminosity also clearly appear in the plot of $RM\Psi_{4}$. For this system, one can also connect the peaks in the gravitational-wave emission to specific events in the binary-single encounter. We again illustrate this connection by adding crosses to the puncture trajectory plots in Figures 11 and 12 at the retarded times $t_{R}=t-R$, which approximately correspond to the moments in time $t$ marked by the dashed lines in Figure 13. One can see that the first peak (colored red) corresponds to the moment when the black holes in the binary get rapidly accelerated out of the orbital plane. This is especially visible in the right plot in Figure 12.
@@
Figure 15: The real part (blue), imaginary part (orange), and the absolute value (green) of $RM\Psi_{4}$ for the system BSS-04, extracted at the radius $R=140M$ and on the level $L=2$ in the direction $\theta=0,\phi=0$ (which corresponds to the $z$-axis).
@@ IV.4 Highly Eccentric Merger After a Strong Binary-Single Scattering (BSS-04)
Our fourth system moves in the $xy$-plane with the normal vector of the binary’s orbital plane pointing in the $z$-direction. For this system we pick $X=50M$, $b=40M$ and $|\vec{p}|=0.057282195M$. The black hole trajectories of the entire system and of the binary in the center-of-mass frame are depicted in Figure 14. Figure 15 shows $RM\Psi_{4}$ in a specific direction ($\theta=0,\phi=0$, which corresponds to the $z$-axis, i.e., the direction of the normal vector of the binary’s orbital plane), extracted at the radius $R=140M$ and on the level $L=2$. One can see that the single black hole induces a significantly accelerated merger with a high eccentricity. In the comparison with the 2.5PN approximation, one can observe that the dynamics look quite similar up to a certain point, although distorted in the direction of motion as in the other examples. The scattering angle appears noticeably larger in the post-Newtonian approximation, which is contrary to what has been observed for the strong-field scattering of individual black holes in [34], where the scattering angles of the higher-order post-Newtonian approximations are always smaller than those of the fully relativistic cases.
@@
Figure 16: Top left: Puncture trajectories of the black holes of BSS-05 (colored blue for the black holes in the binary and orange for the single black hole) with a comparison to the 2.5PN approximation. Top right: Gravitational-wave luminosity extracted at the radius $R=160M$ and on the level $L=2$. Bottom: The real part (blue), the imaginary part (orange), and the absolute value (green) of $RM\Psi_{4}$, extracted at the same radius and level and in the direction $\theta=2.1,\phi=0$.
@@ IV.5 Exchange and Subsequent Merger After a Strong Binary-Single Scattering (BSS-05)
Figure 17: Left: Gravitational-wave luminosity extracted at the radius $R=300M$ and on the level $L=2$. Right: The real part (blue), the imaginary part (orange), and the absolute value (green) of the 22-mode of $RM\Psi_{4}$, extracted at the same radius and level and in the direction $\theta=0.5,\phi=4.9$.
@@
The plot of $RM\Psi_{4}$ in Figure 16 shows the regular binary signal, interrupted by a small bump and a subsequent decay of the signal, which is then followed by a burst of radiation. This can also be seen in the plot of $dE/dt$ where the bump at $t\approx 600M$ is relatively small. Looking at the system at the retarded time $t_{R}=t-R$, one can conclude that the small bump arises from the single black hole slowing down one black hole and accelerating the other one, whereas the burst at the end originates from the merger of the single black hole with one of the binary members. It might be the case that the existence of a bump followed by a binary signal with a different frequency, amplitude, and waveform (or a prompt merger as in this case) is a characteristic feature of an exchange since it requires the deceleration of one binary member’s circular motion and a strong acceleration of the other binary member, leading to a moderate second derivative of the binary’s mass quadrupole moment.
@@ IV.6 Strong Binary-Binary Scattering (BBS-01)
The plot of $dE/dt$ in Figure 17 shows two distinct and irregularly shaped peaks, as well as a smaller constant emission around the time of closest approach (red dashed line). The first peak (colored magenta) likely corresponds to the strong acceleration of both binaries and/or the acceleration of individual binary members. At the time of the first peak, the black holes in the blue-colored binary get accelerated outwards and reach their widest separation throughout the simulation, while at the same time the black-hole distance in the eccentric orange binary reaches a local minimum. The second peak (colored turquoise) is associated with the merger of the orange-colored binary. One can see that the maximum of the absolute value of $RM\Psi_{4}$ occurs earlier than in the gravitational-wave luminosity. This is because the merger sends two bursts in opposite directions. One burst reaches the extraction sphere before the time marked by the turquoise line, while the other one hits it afterward. The plot depicting $RM\Psi_{4}$ also shows the initial signal of both binaries. Due to the different phases, the amplitude gets slightly modulated. The final part of the signal comes from the gravitational-wave emission of the remaining eccentric binary. In the center-of-mass systems of each binary, the black holes get accelerated out of the initial orbital plane during the encounter, and the orbital plane of the binary with blue trajectories gets permanently inclined by around 10 degrees relative to its initial orbital plane.
@@ V Summary and Conclusions
In addition to the previously mentioned difficulties, the extraction of the gravitational-wave signal of binary-single and binary-binary encounters is also challenging due to the fact that the binary is not at rest at the center of the extraction sphere. There are works that deal with the signal extraction for binaries whose center of mass is displaced from the origin or moves at a constant velocity [20, 30, 92], but in our case the binary is accelerating and rotating, which leads to mode mixing and spurious oscillations of $|RM\Psi_{4}|$. When looking at the gravitational-wave luminosity that takes into account all modes, these oscillations are significantly reduced, but some peaks still show irregular shapes, and it is unclear to what extent this is a physical effect or a result of our method of gravitational-wave signal extraction.
@@ Acknowledgments
Animated videos of some of the plots presented in this paper can be found at https://doi.org/10.5281/zenodo.15342415 [55]. They give valuable insights that are not obvious from the plots. The animated version of Figure 9 was created with rose [66]. The code for integrating the three-body post-Newtonian equations of motion up to 2.5PN is available at https://github.com/fmheinze/pn-nbody [56]. All raw data corresponding to the findings in this manuscript are not publicly available upon publication because it is not technically feasible and/or the cost of preparing, depositing, and hosting the data would be prohibitive within the terms of this research project. The data are available from the authors upon reasonable request.
@@
Figure 20: The real part (blue), the imaginary part (orange), and the absolute value (green) of the 22-mode (top panel) and the 44-mode (bottom panel) of $RM\Psi_{4}$ for a quasi-circular black-hole binary moving with a momentum $|\vec{p}|=0.05M$ in the positive $x$-direction, away from the coordinate origin. The signal is extracted at the radius $R=300M$ and on the level $L=3$.
Figure 21: The $l2$ norm of the Hamiltonian constraint (left column) and momentum constraint (right column) of the system BSS-01b (top panels) and BBS-01 (bottom panels) for different resolutions and on two different non-moving levels. For comparison, the gray line indicates the values of a black-hole binary which is fixed at the center of the computational domain. The data is binned in time and the transparent filled regions indicate the total value range for each bin.
Figure 22: Comparison of the Hamiltonian constraint for two binary systems, where one is staggered by half a grid spacing and the other is not staggered.
Appendix A Extra Material
Table 4 provides the full list of parameters used for all the systems in the main text. In Figure 20 we show the 22-mode and the 44-mode of $RM\Psi_{4}$ for a quasi-circular binary that is moving uniformly with respect to the origin of the gravitational-wave extraction sphere as a simple example of mode mixing and the associated spurious oscillations that are discussed in the main text.
Figure 23: The absolute value of $RM\Psi_{4}$ for the system BSS-02 extracted at different radii (indicated with different colors) on the level $L=2$ and in the direction $\theta=2.3,\phi=1.0$. $RM\Psi_{4}$ is reconstructed from all modes from $l=2$ to $l=4$.
Figure 24: Time evolution of the ADM integrals for the mass (first two panels), the $z$-component of the linear momentum (third panel) and the angular momentum (last panel) for the system BSS-01b (solid orange lines) and black-hole binary that is fixed at the center of the grid (dashed gray lines).
Appendix B Numerical Accuracy
In order to test the numerical accuracy of our simulations, we computed the $l2$ norm of the Hamiltonian constraint and the momentum constraint violation for a representative binary-single and binary-binary system. In Figure 21 one can see these quantities plotted and compared with those of a black-hole binary that is fixed at the center of the computational domain (using initial parameters from [88]) and which uses the same numerical setup. For a cleaner visualization, the data is binned in time, and the transparent filled regions indicate the total value range for each bin. The unbinned data look like a baseline (corresponding to the solid lines at the bottom of each shaded region) with several strong spikes that likely occur when a puncture comes close to a grid point. We come to this conclusion because the number of spikes increases with level, resolution, and the number of black holes. Note that we show the norm of the constraints without further normalization, e.g. by second derivatives of the metric, so constraint violations near the punctures are clearly visible. To examine the effect of the distance between punctures and grid points, we show two simulations for the simple binary, where in one case the orbital plane is staggered between the grid points (the default in BAM, as in Figure 21), and in the other case without staggering. The spikes strongly increase when the punctures lie in a plane of grid points and therefore come closer to grid points during the run (see Figure 22). Apart from that, one can see that the constraint violations have a constant to decreasing trend over time. For the Hamiltonian constraint, one can see an improvement with higher resolution, especially when going to $\mathrm{dxyz}=8$. For the momentum constraint, one can only see an improvement with higher resolution for early times. In all cases, the constraint violations for $\mathrm{dxyz}=16$ are comparable or larger by a factor of a few compared to those of the simple binary system. It would be interesting to consider constraint damping as in [39].
In Figure 23 we investigate the consistency of the extracted gravitational-wave signals by plotting $|RM\Psi_{4}|$ for different extraction radii. The relative positions and the approximate shapes of the peaks do not vary much. One can, however, observe that the height of distinct peaks changes differently with radius. This leads to the relative heights of the peaks changing with the extraction radius. Additionally, the spurious oscillations vary slightly for different extraction spheres and they can decrease in amplitude (see e.g. the highest peak in Figure 23). All of this does not significantly change when considering more modes up to $l=6$. Besides the $lm$-modes of $\Psi_{4}$ one also has to consider that under certain conditions, there can be a mixing of the Penrose-Newman scalars $\Psi_{n}$ themselves. The different falloff behaviors of the Penrose-Newman scalars could potentially explain some of the different falloff behaviors observed in our gravitational wave signals. Since the phenomenology and the qualitative aspects of the signals do not change and are consistent across different extraction spheres, and since they all correspond to strong gravitational interactions in the corresponding systems, we consider our statements in the main text to be unaffected by this.
We also tested the conservation of mass, linear momentum, and angular momentum, which is analyzed by monitoring the corresponding ADM surface integrals. This is illustrated in Figure 24 for the binary-single system BSS-01b and the binary-binary system BBS-01, where the relative deviations form the initial values are plotted. These are compared to those of a simple black-hole binary. Initially, one can see strong variations due to junk radiation from imperfect initial data. These variations are the smallest for the binary-binary system. After that, the qualitative behavior of the quantities of the considered systems are very similar. In general, the binary system has the best conservation properties throughout the simulation. The binary-binary system shows slightly worse conservation properties than the binary-single system at late times, especially for the conservation of mass-energy. This is likely due to the more dynamical nature of systems with more black holes, which are not restricted to motion within a plane.